OpenStudy (anonymous):

solve lm+2l +1= 24 can you show work so i know how u got it please ;)

OpenStudy (anonymous):

lm? Do you mean ln?

OpenStudy (anonymous):

\[\left| m+2 \right| +1=24\]

OpenStudy (anonymous):

|m+2|=23 m+2=23 , m+2=-23 so that, m=21 , m=-25

OpenStudy (anonymous):

its 24 not 23

OpenStudy (anonymous):

& thats not the whole problem

OpenStudy (anonymous):

His first step was subtracting the 1 outside the absolute value.

OpenStudy (anonymous):

thanks i understand it now^^^